# SciLitLLM: How to Adapt LLMs for Scientific Literature Understanding

### v3

@@ Abstract
Scientific literature understanding is crucial for extracting targeted information and garnering insights, thereby significantly advancing scientific discovery. Despite the remarkable success of Large Language Models (LLMs), they face challenges in scientific literature understanding, primarily due to (1) a lack of scientific knowledge and (2) unfamiliarity with specialized scientific tasks. To develop an LLM specialized in scientific literature understanding, we propose a hybrid strategy that integrates continual pre-training (CPT) and supervised fine-tuning (SFT), to simultaneously infuse scientific domain knowledge and enhance instruction-following capabilities for domain-specific tasks. In this process, we identify two key challenges: (1) constructing high-quality CPT corpora, and (2) generating diverse SFT instructions. We address these challenges through a meticulous pipeline, including PDF text extraction, parsing content error correction, quality filtering, and synthetic instruction creation. Applying this strategy, we present a suite of LLMs: SciLitLLM, specialized in scientific literature understanding. These models demonstrate promising performance on scientific literature understanding benchmarks. Our contributions are threefold: (1) We present an effective framework that integrates CPT and SFT to adapt LLMs to scientific literature understanding, which can also be easily adapted to other domains. (2) We propose an LLM-based synthesis method to generate diverse and high-quality scientific instructions, resulting in a new instruction set – SciLitIns – for less-represented scientific domains. (3) SciLitLLM achieves promising performance in scientific literature understanding benchmarks. We release the data processing codes [^1] https://github.com/dptech-corp/Uni-SMART/tree/main/SciLitLLM, model weights [^2] https://huggingface.co/Uni-SMART/SciLitLLM-1.5 and the synthetic dataset [^3] https://huggingface.co/datasets/Uni-SMART/SciLitIns.
@@ 1 Introduction
Scientific literature understanding involves the systematic evaluation and interpretation of scientific texts and publications, to identify trends, extract targeted information, and garner insights (AI4Science & Quantum, 2023; Zheng et al., 2023), significantly contributing to scientific discovery. Concurrently, Large Language Models (LLMs) (OpenAI, 2023; Team, 2024; Touvron et al., 2023) have achieved remarkable success in natural language processing, prompting the development of domain-specific LLMs across various fields (Cui et al., 2024; Wu et al., 2023; Clusmann et al., 2023). However, recent studies (Wadden et al., 2024; Cai et al., 2024; Singh et al., 2023) indicate that LLMs face challenges when specializing in scientific literature understanding, particularly in context understanding and question answering. Take Figure 2 as an example, where the LLM is asked to understand the content of a biomedical research paper and then extract the targeted information. LLMs’ potential might be hindered by two major barriers: (1) a lack of scientific knowledge, which results in errors such as the missing important entities in Llama-3.1-8B (AI@Meta, 2024), and (2) unfamiliarity with scientific tasks, leading to the inability of Galactica-6.7B (Taylor et al., 2022) to follow task instructions accurately.
@@
In the CPT stage for domain knowledge injection, we start with an extensive in-house corpus consisting of 73k textbooks and 625k academic papers in the scientific field, all in PDF format. Initially, we leverage PyPDF2444https://pypdf2.readthedocs.io, a widely used open-source PDF parsing tool, to extract raw texts from these documents. We then employ a moderate yet powerful model, Llama3-7B-Instruct, to correct the format and spelling errors introduced by PDF parsing (cf. Section 3.1.1). Subsequently, we train a small text quality classifier to score the corpus and filter out texts of low educational value [^5] Phi models (Gunasekar et al., 2023; Li et al., 2023; Abdin et al., 2024) propose to determine the quality of a pre-training text by its educational value for a student whose goal is to learn basic domain concepts. in the scientific field (cf. Section 3.1.2). These two simple yet effective textual refinement and quality control measures ensure the high quality of our CPT corpus, culminating 12.7 billion tokens for CPT via the Qwen2.5 tokenizer.
@@
Having established such high-quality datasets, we apply the CPT-SFT integration strategy on a general-purpose LLM – Qwen2.5 (Team, 2024) and obtain SciLitLLM of two scales: 7B and 14B. Evaluations on benchmarks of scientific literature understanding demonstrate the effectiveness of our strategy. We observe promising performance enhancements, with an average improvement of 4.0% on SciAssess (Cai et al., 2024) and 10.1% on SciRIFF (Wadden et al., 2024), compared to the leading LLMs under 10B parameters. Notably, SciLitLLM-7B even outperforms Llama3.1 and Qwen2.5 with 70B parameters on SciRIFF. Additionally, SciLitLLM-14B achieves leading results on both benchmarks, surpassing other open-source LLMs. Further ablation studies demonstrate the effectiveness of each module in our pipeline.
@@
2 Related Works
2.1 Knowledge Injection via Continual Pre-training
Pre-training a language model is usually conducted on a large corpus of textual data to learn the statistical properties of language (Radford et al., 2019; Brown et al., 2020). To further inject domain knowledge into a general LLM after pre-training, researchers engage in continual pre-training (CPT) on high-quality domain-specific corpora (Jin et al., 2022; Sun et al., 2020), sometimes combined with general corpora. This process enhances the model’s fundamental understanding abilities in specific downstream domains while mitigating catastrophic forgetting of general knowledge (Wu et al., 2022; Ke et al., 2023; Mehta et al., 2023). See the comprehensive study (Gupta et al., 2023) for different warm-up strategies for CPT. Additionally, the CPT corpora can be augmented by transforming them into an instruction-response format (Cheng et al., 2023; 2024). Furthermore, the scaling law (Hoffmann et al., 2022) of domain-specific CPT (Que et al., 2024) is explored to determine the optimal mix of data. However, these studies primarily focus on training dynamics and data recipes, leaving the pre-processing for scientific data, especially raw PDF files, largely unexplored. Exhibiting such steps is essential for generating high-quality domain corpora and effectively injecting domain knowledge, representing a significant challenge for practitioners.
2.2 Domain Adaptation via Supervised Fine-tuning
Supervised fine-tuning (SFT) modifies a pre-trained language model to follow specific instructions or perform designated tasks by fine-tuning it on a targeted, task-specific dataset (Raffel et al., 2020; Wei et al., 2021). Applying SFT to a general LLM for specific domain adaptation has demonstrated effectiveness in various fields: in medicine (Clusmann et al., 2023), corpora of medical literature and clinical notes are used; in law (Cui et al., 2024), legal documents and case law are compiled; and in finance (Wu et al., 2023), financial reports and market data are utilized. In the scientific domain, several studies have specialized LLMs for scientific tasks, often necessitating the construction of a substantial domain-specific dataset with SFT. For example, SciGLM (Zhang et al., 2024a) leverages existing LLMs to generate step-by-step reasoning for unlabelled scientific instructions. ChemLLM (Zhang et al., 2024b), a more specified LLM in the chemistry field, collects structured chemical data from a vast selection of online databases and transforms this structured data into a question-answering format for SFT. SciRIFF (Wadden et al., 2024) converts existing literature understanding datasets into natural language input-output pairs suitable for instruction-tuning using pre-defined templates. However, benchmark studies (Feng et al., 2024; Cai et al., 2024) indicate that SFT alone may not provide adequate scientific knowledge to excel in relevant tasks. This suggests the need for a more comprehensive approach that combines domain knowledge infusion with instruction-following enhancements.
2.3 LLMs for Scientific Literature Understanding
In the scientific domain, existing strategies for developing specialized LLMs mostly fall into two categories: (1) Supervised fine-tuning with scientific instructions. This approach requires a large, high-quality, and diverse set of instructions to cultivate problem-solving abilities for scientific tasks. Representative works (e.g., SciGLM (Zhang et al., 2024a), ChemLLM (Zhang et al., 2024b), and SciRIFF (Wadden et al., 2024)) have been detailed in Section 2.2. (2) Pre-training with scientific corpora. This approach involves pre-training on a large corpus of scientific texts to improve performance on downstream scientific tasks. Early attempts, such as SciBert (Beltagy et al., 2019) and KV-PLV (Zeng et al., 2022), are based on BERT (Devlin et al., 2019) and pre-trained on a large corpus of scientific text for downstream scientific task enhancement. More recently, Galactica (Taylor et al., 2022) is pre-trained on a vast corpus of scientific literature, including research papers, scientific articles, and other relevant scientific texts. Despite these advances, two major limitations hinder these models from excelling in scientific literature understanding: (1) lack of scientific knowledge, and (2) inability to follow task instructions. To address these challenges, we propose a combined pipeline of CPT and SFT to devise a specialized LLM for scientific literature understanding. It injects domain-specific knowledge through CPT while enhancing task-specific instruction-following abilities through SFT, leading to a more capable LLM for scientific literature understanding.
@@ 3 Method
Figure 5: The pipeline of SciLitLLM consists of two key stages: continual pre-training (CPT) for scientific knowledge injection and supervised fine-tuning (SFT) for scientific instruction following. Specifically, the CPT stage involves several modules: PDF parsing, format & grammar correction (cf. Section 3.1.1), and quality filtering (cf. Section  3.1.2) modules. These modules ensure the model is equipped with high-quality scientific domain knowledge. The SFT stage includes LLM-based instruction generation (cf. Section 3.2.1) and instruction quality control (cf. Section 3.2.2) measures. These steps are designed to fine-tune the model’s ability to follow scientific instructions accurately and effectively.
In this section, we present the details of our proposed pipeline (cf. Figure 5): continual pre-training for scientific knowledge injection and supervised fine-tuning for scientific tasks enhancement.
@@ 3.1 CPT for Scientific Knowledge Injection
3.1.1 Format & Grammar Correction
A parsed text from a PDF document often contains many formatting and syntax errors. To address this issue, we prompt Llama3-8B-Instruct, to correct these errors introduced during the PDF parsing process. Utilizing the vLLM (Kwon et al., 2023) backend, Llama3-8B-Instruct can process approximately 2.52 million tokens per Nvidia A100 GPU hour. The process takes over 5,000 A100 GPU hours to handle all the textbooks and research papers. Example texts – both before and after processing – along with the prompt template are provided in Appendix A to demonstrate the improvements through this correction process.
3.1.2 CPT Quality Filter
During CPT, maintaining the quality of the training corpus is crucial for effective knowledge injection. Given the extremely large scale of pre-training corpora, assessing quality through human annotation is not feasible (Erdmann et al., 2019; Qiu et al., 2023). Consequently, leading LLMs (e.g., Phi (Gunasekar et al., 2023), Llama (Touvron et al., 2023), and Qwen2.5 (Team, 2024)) employ model-based quality filters. The typical process involves using larger LLMs to score the quality of a subset of texts, which then serve as labels for training small classifiers (e.g., random forest (Gunasekar et al., 2023) and Bert (AI@Meta, 2024)) to annotate the entire training corpus. Inspired by this approach, we design a resource-efficient method based on a lightweight text quality classifier.
This process results in a scientific text quality classifier tailored for scientific corpus assessment. We then utilize this classifier to assess the quality of the entire CPT dataset and exclude the lowest-scoring 25% (See Figure 4 for concrete samples). See Appendix B for more details about classifier training and hyperparameters.
3.1.3 CPT Training Settings
@@
After preparing all the corpus, we perform CPT on Qwen2.5-Base (Team, 2024) for one epoch, encompassing 23.7 billion tokens (cf. Table 1), with a sequence length of 2,048 tokens. To maintain the model’s general knowledge, we also include a similar scale of general corpus tokens from Redpajama (Computer, 2023). To stabilize the learning procedure, we gradually decrease the learning rate from $1\times 10^{-5}$ to 0 with a cosine scheduler. To address overfitting, we apply a weight decay of $0.1$ and gradients were clipped at a maximum value of $1.0$. The CPT training took approximately 3 days on 32 Nvidia A100 GPUs for SciLitLLM-7B-Base and about 7 days for the 14B model.
@@ 3.2 SFT for Scientific Instruction Following
  • Existing instruction-tuning datasets in the scientific domain (Feng et al., 2024; Fang et al., 2023; Li et al., 2024) primarily focus on fields such as physics, chemistry, and biology. Manually collecting instruction-tuning data for other less-represented vertical domains (e.g., biomedicine, and material) is both time-consuming and costly (Erdmann et al., 2019; Qiu et al., 2023).
@@ 3.2.1 Instruction Synthesis of Less-represented Domains
Utilizing this pipeline, we obtain over 100k synthetic instructions covering less-represented scientific domains and various types of specialized tasks (details are presented in Appendix C.1).
@@ 3.2.2 Instruction Quality Control
    Heuristic deduplication. Despite the measures taken during the generation process to prevent high homogeneity in the instructions, the generated data points may still contain similar questions or identical answers. To eliminate such redundancy, we implement a deduplication process to remove similar data points. See Appendix C.2 for details.
@@
    LLM-based filtering. Inspired by recent efforts (Eldan & Li, 2023; Chiang & Lee, 2023; Zhang et al., 2024c) to measure the quality of generated content using LLMs, we leverage Llama-3-70B-Instruct to assess the quality of generated instructions for five aspects: clarity, complexity, correctness, usefulness, and adaptability. We show the quality statistics of synthetic instructions in Figure 7. The detailed recipe for instruction quality evaluation with concrete examples is included in Appendix C.3.
@@ 4 Experiments
Llama3.1-8B 56.35 25.88 44.50 53.28 45.00
Qwen2.5-7B 68.36 48.37 51.64 56.23 56.15
SciLitLLM-7B 70.45 51.21 54.06 59.27 58.75
@@
Qwen2.5-14B 79.78 62.73 62.22 66.92 67.91
SciLitLLM-14B 80.62 64.44 64.24 68.06 69.34
Table 2: Performance comparison of base models. SciLitLLM outperforms Llama3.1 and Qwen2.5 with similar scales, demonstrating improved domain-specific understanding through continual pretraining on scientific corpora.
@@
SciTulu-7B Mistral-7B Llama3.1-8B Qwen2.5-7B SciLitLLM-7B Qwen2.5-14B Llama3.1-70B Qwen2.5-72B SciLitLLM-14B GPT3.5 GPT4o
SciAssess Biology 45.3 52.0 63.4 60.9 65.3 65.9 69.4 67.2 67.0 55.4 68.9
Chemistry 19.0 34.0 46.1 47.9 55.4 60.0 62.5 63.7 63.4 37.3 68.0
Material 31.3 36.5 50.9 48.9 53.7 57.9 59.2 59.4 62.6 37.0 62.0
Medicine 19.9 25.1 30.2 28.3 32.4 31.5 37.4 37.8 39.9 31.8 45.8
@@
SciRIFF BioASQ 37.5 43.2 45.0 45.3 51.2 48.0 43.5 46.7 53.9 47.3 46.7
BioR 55.7 48.9 48.1 43.9 76.6 56.8 61.5 57.8 83.3 53.9 61.0
DiscMT 61.5 44.5 68.8 73.7 71.1 70.5 77.3 54.7 77.7 67.9 78.3
EI 11.6 17.2 17.2 23.3 23.5 27.2 26.0 26.3 30.8 19.2 24.7
MC 34.6 47.5 49.7 50.8 70.7 54.2 60.9 58.0 72.1 47.8 58.7
MuP 72.1 87.5 87.0 84.9 67.5 91.5 90.9 94.0 67.7 76.8 86.9
Qasper 54.2/38.6 48.8/43.2 46.5/42.7 38.1/25.7 50.7/54.1 48.4/46.8 49.1/40.2 61.9/50.5 54.0/55.1 54.7/39.8 67.8/50.5
SciERC 35.6 31.1 32.9 30.1 49.9 33.7 36.2 34.1 54.7 28.6 42.2
SciFact 66.0/49.2 68.8/53.8 63.8/53.7 77.8/60.3 83.5/67.3 83.1/65.6 87.3/70.3 86.5/69.6 91.0/68.6 69.7/53.3 84.3/68.7
@@
Table 3: Model performances on scientific literature understanding benchmarks: SciAssess and SciRIFF. SciLitLLM-7B and SciLitLLM-14B achieve leading performance compared with open-source models of similar scales. The best-performing models smaller than 10B and larger than 10B are highlighted in bold. Results for SciTulu-7B, GPT-3.5, and GPT-4o on SciRIFF are taken from its original papers, while all other results are produced by us.

### v4

@@ Abstract
Scientific literature understanding is crucial for extracting targeted information and garnering insights, thereby significantly advancing scientific discovery. Despite the remarkable success of Large Language Models (LLMs), they face challenges in scientific literature understanding, primarily due to (1) a lack of scientific knowledge and (2) unfamiliarity with specialized scientific tasks. To develop an LLM specialized in scientific literature understanding, we propose a hybrid strategy that integrates continual pre-training (CPT) and supervised fine-tuning (SFT), to simultaneously infuse scientific domain knowledge and enhance instruction-following capabilities for domain-specific tasks. In this process, we identify two key challenges: (1) constructing high-quality CPT corpora, and (2) generating diverse SFT instructions. We address these challenges through a meticulous pipeline, including PDF text extraction, parsing content error correction, quality filtering, and synthetic instruction creation. Applying this strategy, we present a suite of LLMs: SciLitLLM, specialized in scientific literature understanding. These models demonstrate promising performance on scientific literature understanding benchmarks. Our contributions are threefold: (1) We present an effective framework that integrates CPT and SFT to adapt LLMs to scientific literature understanding, which can also be easily adapted to other domains. (2) We propose an LLM-based synthesis method to generate diverse and high-quality scientific instructions, resulting in a new instruction set – SciLitIns – for less-represented scientific domains. (3) SciLitLLM achieves promising performance in scientific literature understanding benchmarks. We release the data processing codes [^1] https://github.com/dptech-corp/Uni-SMART and model weights [^2] https://huggingface.co/collections/Uni-SMART/scilitllm15-67283353ada975ba995629ef.
@@ 1 Introduction
Scientific literature understanding involves the systematic evaluation and interpretation of scientific texts and publications, to identify trends, extract targeted information, and garner insights (AI4Science & Quantum, 2023; Zheng et al., 2023), significantly contributing to scientific discovery. Concurrently, Large Language Models (LLMs) (OpenAI, 2023; Qwen, 2024; Touvron et al., 2023) have achieved remarkable success in natural language processing, prompting the development of domain-specific LLMs across various fields (Cui et al., 2024; Wu et al., 2023; Clusmann et al., 2023). However, recent studies (Wadden et al., 2024; Cai et al., 2024; Singh et al., 2023) indicate that LLMs face challenges when specializing in scientific literature understanding, particularly in context understanding and question answering. Take Figure 2 as an example, where the LLM is asked to understand the content of a biomedical research paper and then extract the targeted information. LLMs’ potential might be hindered by two major barriers: (1) a lack of scientific knowledge, which results in errors such as the missing important entities in Llama-3.1-8B (AI@Meta, 2024), and (2) unfamiliarity with scientific tasks, leading to the inability of Galactica-6.7B (Taylor et al., 2022) to follow task instructions accurately.
@@
In the CPT stage for domain knowledge injection, we start with an extensive in-house corpus consisting of 73k textbooks and 625k academic papers in the scientific field, all in PDF format. Initially, we leverage PyPDF2333https://pypdf2.readthedocs.io, a widely used open-source PDF parsing tool, to extract raw texts from these documents. We then employ a moderate yet powerful model, Llama3-8B-Instruct, to correct the format and spelling errors introduced by PDF parsing (cf. Section 2.1.1). Subsequently, we train a small text quality classifier to score the corpus and filter out texts of low educational value [^4] Phi models (Gunasekar et al., 2023; Li et al., 2023; Abdin et al., 2024) propose to determine the quality of a pre-training text by its educational value for a student whose goal is to learn basic domain concepts. in the scientific field (cf. Section 2.1.2). These two simple yet effective textual refinement and quality control measures ensure the high quality of our CPT corpus, culminating 12.7 billion tokens for CPT via the Qwen2.5 tokenizer.
@@
Having established such high-quality datasets, we apply the CPT-SFT integration strategy on a general-purpose LLM – Qwen2.5 (Qwen, 2024) and obtain SciLitLLM of two scales: 7B and 14B. Evaluations on benchmarks of scientific literature understanding demonstrate the effectiveness of our strategy. We observe promising performance enhancements, with an average improvement of 4.0% on SciAssess (Cai et al., 2024) and 10.1% on SciRIFF (Wadden et al., 2024), compared to the leading LLMs under 10B parameters. Notably, SciLitLLM-7B even outperforms Llama3.1 and Qwen2.5 with 70B parameters on SciRIFF. Additionally, SciLitLLM-14B achieves leading results on both benchmarks, surpassing other open-source LLMs. Further ablation studies demonstrate the effectiveness of each module in our pipeline.
@@ 2 Method
Figure 5: The pipeline of SciLitLLM consists of two key stages: continual pre-training (CPT) for scientific knowledge injection and supervised fine-tuning (SFT) for scientific instruction following. Specifically, the CPT stage involves several modules: PDF parsing, format & grammar correction (cf. Section 2.1.1), and quality filtering (cf. Section  2.1.2) modules. These modules ensure the model is equipped with high-quality scientific domain knowledge. The SFT stage includes LLM-based instruction generation (cf. Section 2.2.1) and instruction quality control (cf. Section 2.2.2) measures. These steps are designed to fine-tune the model’s ability to follow scientific instructions accurately and effectively.
In this section, we present the details of our proposed pipeline (cf. Figure 5): continual pre-training for scientific knowledge injection and supervised fine-tuning for scientific tasks enhancement. We refer to Appendix A for a comprehensive literature review on knowledge injection, domain adaptation, and scientific literature understanding for LLMs.
@@ 2.1 CPT for Scientific Knowledge Injection
2.1.1 Format & Grammar Correction
A parsed text from a PDF document often contains many formatting and syntax errors. To address this issue, we prompt Llama3-8B-Instruct, to correct these errors introduced during the PDF parsing process. Utilizing the vLLM (Kwon et al., 2023) backend, Llama3-8B-Instruct can process approximately 2.52 million tokens per Nvidia A100 GPU hour. The process takes over 5,000 A100 GPU hours to handle all the textbooks and research papers. Example texts – both before and after processing – along with the prompt template are provided in Appendix B to demonstrate the improvements through this correction process.
2.1.2 CPT Quality Filter
Figure 6: Score distribution of the CPT Data. The filtered-out 25% data are marked gray, while the remaining 75% are marked orange.
Figure 7: The quality of SciLitIns is evaluated from five aspects: clarity, complexity, correctness, usefulness, and adaptability. Instructions with an average score of less than 4 are filtered out.
During CPT, maintaining the quality of the training corpus is crucial for effective knowledge injection. Given the extremely large scale of pre-training corpora, assessing quality through human annotation is not feasible (Erdmann et al., 2019; Qiu et al., 2023). Consequently, leading LLMs (e.g., Phi (Gunasekar et al., 2023), Llama (Touvron et al., 2023), and Qwen2.5 (Qwen, 2024)) employ model-based quality filters. The typical process involves using larger LLMs to score the quality of a subset of texts, which then serve as labels for training small classifiers (e.g., random forest (Gunasekar et al., 2023) and Bert (AI@Meta, 2024)) to annotate the entire training corpus. Inspired by this approach, we design a resource-efficient method based on a lightweight text quality classifier.
This process results in a scientific text quality classifier tailored for scientific corpus assessment. We then utilize this classifier to assess the quality of the entire CPT dataset and exclude the lowest-scoring 25% (See Figure 4 for concrete samples). The distribution of the scores is illustrated in Figure 7. We refer to Appendix C for more details about classifier training and sensitivity analysis of the filtering threshold.
2.1.3 CPT Training Settings
After preparing all the corpus, we perform CPT on Qwen2.5-Base (Qwen, 2024) for one epoch, encompassing 23.7 billion tokens (cf. Table 1), with a sequence length of 2,048 tokens. To maintain the model’s general knowledge, we also include a similar scale of general corpus tokens from Redpajama (Computer, 2023). To stabilize the learning procedure, we gradually decrease the learning rate from $1\times 10^{-5}$ to 0 with a cosine scheduler. To address overfitting, we apply a weight decay of $0.1$ and gradients were clipped at a maximum value of $1.0$. The CPT training took approximately 3 days on 32 Nvidia A100 GPUs for SciLitLLM-7B-Base and about 7 days for the 14B model.
@@ 2.2 SFT for Scientific Instruction Following
  • Existing instruction-tuning datasets in the scientific domain (Feng et al., 2024; Fang et al., 2023) primarily focus on fields such as physics, chemistry, and biology. Manually collecting instruction-tuning data for other less-represented vertical domains (e.g., biomedicine, and material) is both time-consuming and costly (Erdmann et al., 2019; Qiu et al., 2023).
@@ 2.2.1 Instruction Synthesis of Less-represented Domains
Utilizing this pipeline, we obtain over 100k synthetic instructions covering less-represented scientific domains and various types of specialized tasks (details are presented in Appendix D.1).
@@ 2.2.2 Instruction Quality Control
    Heuristic deduplication. Despite the measures taken during the generation process to prevent high homogeneity in the instructions, the generated data points may still contain similar questions or identical answers. To eliminate such redundancy, we implement a deduplication process to remove similar data points. See Appendix D.2 for details.
@@
    LLM-based filtering. Inspired by recent efforts (Eldan & Li, 2023; Chiang & Lee, 2023; Zhang et al., 2024c) to measure the quality of generated content using LLMs, we leverage Llama-3-70B-Instruct to assess the quality of generated instructions for five aspects: clarity, complexity, correctness, usefulness, and adaptability. We show the quality statistics of synthetic instructions in Figure 7. Instructions with an average score of less than 4 are filtered out. The detailed recipe for instruction quality evaluation with concrete examples is included in Appendix D.3.
@@ 3 Experiments
Llama3.1-8B-Base 56.35 25.88 44.50 53.28 45.00
Qwen2.5-7B-Base 68.36 48.37 51.64 56.23 56.15
SciLitLLM-7B-Base 70.45 51.21 54.06 59.27 58.75
@@
Qwen2.5-14B-Base 79.78 62.73 62.22 66.92 67.91
SciLitLLM-14B-Base 80.62 64.44 64.24 68.06 69.34
Table 2: Accuracy(%) comparison of base models. SciLitLLM outperforms Llama3.1 and Qwen2.5 with similar scales, demonstrating improved domain-specific understanding through continual pretraining on scientific corpora. Underline indicates those trained on domain-specific corpora.
@@
Mistral-7B Llama3.1-8B Qwen2.5-7B SciTulu-7B SciLitLLM-7B Qwen2.5-14B Llama3.1-70B Qwen2.5-72B SciLitLLM-14B GPT3.5 GPT4o
SciAssess Biology 52.0 63.4 60.9 45.3 65.3 65.9 69.4 67.2 67.0 55.4 68.9
Chemistry 34.0 46.1 47.9 19.0 55.4 60.0 62.5 63.7 63.4 37.3 68.0
Material 36.5 50.9 48.9 31.3 53.7 57.9 59.2 59.4 62.6 37.0 62.0
Medicine 25.1 30.2 28.3 19.9 32.4 31.5 37.4 37.8 39.9 31.8 45.8
@@
SciRIFF BioASQ 43.2 45.0 45.3 37.5 51.2 48.0 43.5 46.7 53.9 47.3 46.7
BioR 48.9 48.1 43.9 55.7 76.6 56.8 61.5 57.8 83.3 53.9 61.0
DiscMT 44.5 68.8 73.7 61.5 71.1 70.5 77.3 54.7 77.7 67.9 78.3
EI 17.2 17.2 23.3 11.6 23.5 27.2 26.0 26.3 30.8 19.2 24.7
MC 47.5 49.7 50.8 34.6 70.7 54.2 60.9 58.0 72.1 47.8 58.7
MuP 87.5 87.0 84.9 72.1 67.5 91.5 90.9 94.0 67.7 76.8 86.9
Qasper 48.8/43.2 46.5/42.7 38.1/25.7 54.2/38.6 50.7/54.1 48.4/46.8 49.1/40.2 61.9/50.5 54.0/55.1 54.7/39.8 67.8/50.5
SciERC 31.1 32.9 30.1 35.6 49.9 33.7 36.2 34.1 54.7 28.6 42.2
SciFact 68.8/53.8 63.8/53.7 77.8/60.3 66.0/49.2 83.5/67.3 83.1/65.6 87.3/70.3 86.5/69.6 91.0/68.6 69.7/53.3 84.3/68.7
@@
Table 3: Model performances on scientific literature understanding benchmarks: SciAssess and SciRIFF. SciLitLLM-7B and SciLitLLM-14B achieve leading performance compared with open-source models of similar scales. The best-performing models smaller than 10B and larger than 10B are highlighted in bold. Results for SciTulu-7B, GPT-3.5, and GPT-4o on SciRIFF are taken from its original papers, while all other results are produced by us. Underline indicates those trained on domain-specific instructions.